# SoulX-FlashHead: Oracle-guided Generation of Infinite Real-time Streaming Talking Heads

### v1

@@ 1 Introduction
SoulX-FlashHead is provided in two flexible deployment versions. SoulX-FlashHead-Lite targets ultra-fast interaction scenarios and achieves real-time inference on a single RTX 4090 GPU. SoulX-FlashHead-Base pursues superior detail and enables real-time generation on dual RTX 5090 GPUs.
@@ 2.1.2 Data filter and annotation stage
In the data filtering stage, we utilize the SyncNet (Chung and Zisserman, 2016) model to calculate LSE-C and LSE-D confidence scores and discard samples with poor alignment. To ensure consistent facial visibility, we inspect raw footage at a sampling rate of one frame per second and exclude videos where the proportion of frames lacking a detectable face exceeds a predefined threshold. We further employ DWpose (Yang et al., 2023) to extract body keypoints and strictly filter out clips exhibiting hand-over-face occlusion to prevent generation artifacts. Optical flow (Dosovitskiy et al., 2015) analysis serves to identify and remove sequences containing scene discontinuities. Valid clips are finally cropped to a resolution of $512\times 512$ pixels centered on the detected face.
@@ 3.1 Model Architecture
3D-VAE. To strictly address the trade-off between visual fidelity and inference latency, we employ distinct VAE architectures for our Base and Lite variants. The Base model utilizes the WAN 2.1 VAE (Wan et al., 2025) which encodes video frames into compact latent representations with a spatio-temporal downsampling factor of $4\times 8\times 8$. This configuration prioritizes high-fidelity reconstruction and fine-grained detail preservation suitable for quality-critical scenarios. Conversely, the Lite model incorporates the LTX-VAE (HaCohen et al., 2024) to optimize for real-time performance. LTX-VAE achieves aggressive compression by downsampling inputs by a factor of $(32,32,8)$, yielding a pixel-to-token ratio of 8192:18192:1—approximately $32\times$ higher than the WAN scheme. While LTX-VAE integrates a diffusion decoding step to mitigate detail loss inherent to such high compression, it fundamentally serves as the engine for speed-balanced inference in latency-sensitive applications.
@@ 4.1 Performance of SoulX-FlashHead
SoulX-FlashHead (Base) 13.67 133.69 5.60 7.81 -
Main Results. We performed a comprehensive quantitative comparison with state-of-the-art (SOTA) methods on two benchmark datasets: HDTF (Zhang et al., 2021) and VFHQ (Xie et al., 2022). The quantitative results are summarized in Tab. 3 and Tab. 4. On the HDTF dataset, the non-streaming Base model achieves an FID of 8.31 and outperforms EchoMimic. In the streaming setting, the Base model yields an FID of 9.97 which surpasses diffusion baselines including Hallo3 and AniPortrait. Regarding video smoothness measured by FVD, the Base model scores 103.14 in non-streaming and 111.38 in streaming modes. Both results improve upon Sonic which scores 113.31, indicating that our method generates videos with high temporal coherence. Experiments on the VFHQ dataset demonstrate significant advantages in lip synchronization. The Base model achieves Sync-C scores of 5.60 for non-streaming and 5.53 for streaming. These scores exceed both SadTalker at 4.49 and Sonic at 4.64. These results validate the effectiveness of the Oracle-Guided Bidirectional Distillation strategy in aligning lip movements with audio signals and maintaining synchronization precision in complex natural scenarios. Regarding inference efficiency and streaming strategy analysis, the Lite model achieves an inference speed of 96 FPS. Compared to the lightweight method Ditto, the Lite model maintains real-time performance while delivering superior visual quality. Against computationally intensive models such as Hallo3 and EchoMimic, our approach offers a substantial speed advantage. A comparison between streaming and non-streaming data reveals minimal performance degradation. For instance, on HDTF, the streaming Lite model retains a Sync-C score of 4.21 and a low FVD. This confirms that the Temporal Audio Context Cache effectively mitigates audio feature collapse during real-time streaming while the motion frame sampling strategy successfully alleviates cold-start issues.
In conclusion, the Base model achieves state-of-the-art visual and synchronization quality, while the Lite model strikes an optimal balance between high fidelity and real-time responsiveness, outperforming existing methods in comprehensive capability.
@@ 5 Conclusion
This work introduces SoulX-FlashHead which is a unified framework that effectively reconciles the conflict between high-fidelity video generation and low-latency streaming interaction. By leveraging a 1.3B-parameter DiT backbone, we implemented a robust two-stage training strategy. The Streaming-Aware Spatiotemporal Pre-training ensures stable audio-visual alignment under unstable streaming conditions while the Oracle-Guided Bidirectional Distillation significantly mitigates identity drift and error accumulation in long-sequence autoregressive generation. Extensive experiments demonstrate that our approach achieves state-of-the-art performance on benchmark datasets and offers flexible deployment options ranging from the ultra-fast 96 FPS Lite version to the high-fidelity Base version.

### v2

@@ 1 Introduction
SoulX-FlashHead is provided in two flexible deployment versions. SoulX-FlashHead-Lite targets ultra-fast interaction scenarios and achieves real-time inference on a single RTX 4090 GPU. SoulX-FlashHead-Pro pursues superior detail and enables real-time generation on dual RTX 5090 GPUs.
@@ 2.1.2 Data filter and annotation stage
In the data filtering phase, we first perform face detection and mask extraction on the initial frame of each clip, and valid clips are finally cropped to a resolution of $512\times 512$ pixels centered on the detected face. Subsequently, optical flow analysis (Dosovitskiy et al., 2015) is employed to identify and remove sequences containing abrupt scene transitions. To ensure consistent facial visibility, we inspect raw footage at a sampling rate of one frame per second and exclude videos where the proportion of frames lacking a detectable face exceeds a predefined threshold. We further utilize DWpose (Yang et al., 2023) to extract body keypoints and strictly filter out clips exhibiting hand-over-face occlusion to prevent generation artifacts. Finally, we employ the SyncNet (Chung and Zisserman, 2016) model to calculate LSE-C and LSE-D confidence scores to discard samples with poor audio-visual alignment.
@@ 3.1 Model Architecture
3D-VAE. To strictly address the trade-off between visual fidelity and inference latency, we employ distinct VAE architectures for our Pro and Lite variants. The Pro model utilizes the WAN 2.1 VAE (Wan et al., 2025) which encodes video frames into compact latent representations with a spatio-temporal downsampling factor of $4\times 8\times 8$. This configuration prioritizes high-fidelity reconstruction and fine-grained detail preservation suitable for quality-critical scenarios. Conversely, the Lite model incorporates the LTX-VAE (HaCohen et al., 2024) to optimize for real-time performance. LTX-VAE achieves aggressive compression by downsampling inputs by a factor of $(32,32,8)$, yielding a pixel-to-token ratio of 8192:18192:1—approximately $32\times$ higher than the WAN scheme. While LTX-VAE integrates a diffusion decoding step to mitigate detail loss inherent to such high compression, it fundamentally serves as the engine for speed-balanced inference in latency-sensitive applications.
@@ 4.1 Performance of SoulX-FlashHead
SoulX-FlashHead (Pro) 13.67 133.69 5.60 7.81 -
Main Results. We performed a comprehensive quantitative comparison with state-of-the-art (SOTA) methods on two benchmark datasets: HDTF (Zhang et al., 2021) and VFHQ (Xie et al., 2022). The quantitative results are summarized in Tab. 3 and Tab. 4. On the HDTF dataset, the non-streaming Pro model achieves an FID of 8.31 and outperforms EchoMimic. In the streaming setting, the Pro model yields an FID of 9.97 which surpasses diffusion baselines including Hallo3 and AniPortrait. Regarding video smoothness measured by FVD, the Pro model scores 103.14 in non-streaming and 111.38 in streaming modes. Both results improve upon Sonic which scores 113.31, indicating that our method generates videos with high temporal coherence. Experiments on the VFHQ dataset demonstrate significant advantages in lip synchronization. The Pro model achieves Sync-C scores of 5.60 for non-streaming and 5.53 for streaming. These scores exceed both SadTalker at 4.49 and Sonic at 4.64. These results validate the effectiveness of the Oracle-Guided Bidirectional Distillation strategy in aligning lip movements with audio signals and maintaining synchronization precision in complex natural scenarios. Regarding inference efficiency and streaming strategy analysis, the Lite model achieves an inference speed of 96 FPS. Compared to the lightweight method Ditto, the Lite model maintains real-time performance while delivering superior visual quality. Against computationally intensive models such as Hallo3 and EchoMimic, our approach offers a substantial speed advantage. A comparison between streaming and non-streaming data reveals minimal performance degradation. For instance, on HDTF, the streaming Lite model retains a Sync-C score of 4.21 and a low FVD. This confirms that the Temporal Audio Context Cache effectively mitigates audio feature collapse during real-time streaming while the motion frame sampling strategy successfully alleviates cold-start issues.
In conclusion, the Pro model achieves state-of-the-art visual and synchronization quality, while the Lite model strikes an optimal balance between high fidelity and real-time responsiveness, outperforming existing methods in comprehensive capability.
@@ 5 Conclusion
This work introduces SoulX-FlashHead which is a unified framework that effectively reconciles the conflict between high-fidelity video generation and low-latency streaming interaction. By leveraging a 1.3B-parameter DiT backbone, we implemented a robust two-stage training strategy. The Streaming-Aware Spatiotemporal Pre-training ensures stable audio-visual alignment under unstable streaming conditions while the Oracle-Guided Bidirectional Distillation significantly mitigates identity drift and error accumulation in long-sequence autoregressive generation. Extensive experiments demonstrate that our approach achieves state-of-the-art performance on benchmark datasets and offers flexible deployment options ranging from the ultra-fast 96 FPS Lite version to the high-fidelity Pro version.